In a certain school, there are 456 girls. Calculate the total number of students if 24% of the total students are boys.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Letthe total number of students to be 100.

It is given that 24% are boys.

So, total number of boys = 24% of 100 = 24

Thus, the total number of girls will be = (100-24)% i.e. 76%.

So, the total number of girls = 76% of 100 = 76

But, it is given that there are 456 girls.

Now, for 76 girls, total students are 100

For, 1 girl, total students will be 100/76

Therefore, for 456 girls, the total number of students will be = 100/76 × 456 = 45600/76 = 600

Hence, the total number of students = 600.
Thanks